OpenStudy (anonymous):

Calculate the discriminant. x2 – 2x + 2 = 0 -__- I have no idea what discriminant is?

OpenStudy (diyadiya):

\[Discriminant = b^2 - 4ac\]

OpenStudy (diyadiya):

Here a=1 b= -2 c=2

OpenStudy (apoorvk):

When you compare a quadratic equation with the standard form, i.e., \(ax^2+bx+c=0\) Then the discriminant 'D' is: \[D = b^2 - 4ac\]

OpenStudy (anonymous):

D= -4

OpenStudy (anonymous):

so it's -2 +/- \[\sqrt{4-8}\]

OpenStudy (anonymous):

\[-2\pm \sqrt{4-8}\]

OpenStudy (diyadiya):

there's no radical its just \(b^2 - 4ac \)

OpenStudy (anonymous):

oh. so the two answers are -6 and 4

OpenStudy (diyadiya):

How?

OpenStudy (diyadiya):

can you show?

OpenStudy (anonymous):

well LOL I did the plus and minus? I dunno that's how I remember the quadratic formula after you solve everything? i'm stupid so LOL

OpenStudy (anonymous):

-2 - -6 and -2 + -6

OpenStudy (diyadiya):

No, They just asked to calculate Discriminant You don't have to use the entire quadratic formula here \[Discriminant = b^2 - 4ac = (-2)^2 - 4*1 * 2 = 4 - 8= ?\]

OpenStudy (anonymous):

-4 ?

OpenStudy (diyadiya):

Right! :)